Eliminate the fractions (write all the coefficients and the constant as integers) in the following equation: x(3/4) + y(2/3) + z(5/6) = 2.
Answer by mananth(16949)

x(3/4) + y(2/3) + z(5/6) = 2.
...
LCM= 12
...
9x+8y+10z / 12=2
9x+8y+10z=12*2
9x+8y+10z=24